OpenStudy (anonymous):

Consider the following right triangle. Side a has length 5 cm and side c has length 2√10 cm. Find the length of side b.

OpenStudy (callisto):

Which one has the longest side? a, b, or c?

OpenStudy (callisto):

Use Pyth. Thm. \[a^2 + b^b = c^2\] Pur a=5, c= 2√10 into the above formula \[5^2 + b^2 = (2\sqrt{10})^2\] \[ b^2 = 40 -15\] Can you continue from here?

OpenStudy (callisto):

Sorry, it should be \[b^2 = 40-25\] Typing mistake

OpenStudy (callisto):

b would be equal to √15

OpenStudy (anonymous):

ok im back and i didnt even get A notification tellling me you wrote on here sorry

OpenStudy (callisto):

No worry, It's okay :) Do you understand my workings?

OpenStudy (anonymous):

Yes i do thank you for working out most people give me the answe and i hate that if im going to ask for help i want help working out the problem and not just the answer so thank you so so so much :)

OpenStudy (callisto):

Welcome :)

OpenStudy (anonymous):

:))